Loading...
ST8402-CS 890927Date: To: From: Subject: MEMORANDUM September 27, 1989 John C. Karlsruher, P.E. Weekly Progress Report on Denton Tap Road Projects Tuesday, 9/19/89 (continuation of last week's report) South project - L.H. Lacy Co. - clean-up crew #2, i foreman, 2 blades, 4 laborers, I loader. Topsoil in median islands and east outside back of curbs. Jim Bowman Const. Co. - James Long, 3 men. Tieing rail steel for east 1/2 of bridge and bridge clean- up. North project - Lectric Safety Lites - did not come back to project to install 8" yellow buttons, as previously stated. Note: traffic switch scheduled for tomorrow .... Wednesday, 9/20/S9 Clear, Sunny 60°-86° North project - L.H. Lacy Co. - Clean-up crew %1, Steve - 1 loader, 3 to'4 laborers. Power broom sweeping east 34' width pavement. Lectric Safety Lites - Greg Morrison - Installing barricades for lane tapers, in lieu of pavement parkings and striping. Button crew did not show,, again, to install 8" yellow buttons and stripes and the re-working of the left turn lane buttons prior to opening of the new roadway. Brooks Maltby - 3 men, 1 backhoe. Mobilized on project at 10:00 AM ± to install pull boxes and PVC sweeps. Started at Parkway Blvd./Denton Tap Road intersection. Eagle Const. Co. - Greg, 4 men. Working grade within median nose sta. 170+64 for schedule median nose cap concrete placement at 1:30 PM. Re-worked damaged redwood header and dowel bars. Time: 11:30 AM - 3:30 PM. Arjang Electric (Traffic signal Contractor) - 2 men, bucket truck, with Rick Campbell at 11:00 AM. Repositioning signal heads for L.H. Lacy's traffic switch. ' Time: 9:00 AM - 4:00 PM. Note: switched traffic at 3:35 PM. Signals operational and programmed. SYB Const. Co. - Randy Bennett, 2 men. Cutting 14" casing pipe around 8" PVC main to Skaggs store from 2:45 PM - 6:30 PM. city hooking up domestic by-pass for Skaggs, since isolation of section to be worked on shut off their water. As-built drawings for shopping center reflect isolation of pipe within this section would not have any business out of water, city hooked up by-pass from 3:00 PM - 5:00 PM. Tommy Estes Co. - Did not show on project as stated yesterday, for printing of the 4" white stripe at the north transition, right centerline sta. 191+75 to 206+30 ±. L.H. Lacy Co. - 4:30 PM - Asst. Proj. Manager informed me that the button and striping crews would not be back on project until Monday 9/25/89. Note: L.H. Lacy Co. did not submit a barricade layout submittal for the traffic switch. Project Manager departed site at 2:00 PM and did not inform anyone on the means of handling traffic, etc. Discussed with Asst Project. Manager (Ross) and he did not know what was going on. Asked barricade techn~¢lan and he replied, don't have any information or drawings on the proposed layout and how L.H. Lacy intends to handle traffic". South project - L.H. Lacy Co. - 2nd clean-up crew, 1 loader, box blade, 1 blade, 5 laborers. #2 blade was broken down most of the day. (Parked in middle of east 34' width roadway right ' ' n ' centerline sta. 142+00). Removed HMAC trans~t~o across 3 width median island at Sandy Lake - Denton Tap Road. Placing topsoil in median islands. Weedeater being used to trim 4' high weeds in west parkway along Mr. Thweatt's property 149+00 to 140+00 ±. Had a scheduled meeting for 1:00 PM today with Ron Van Ness and Steve Parsons and Scott Balch (Project Manager for L.H. Lacy). Scott Balch did not show up for the meeting.i Discussed issues with the property owners until 1:45 PM. Jim Bowman Const. Co. - Did not show - no work on bridge rails. Eagle Const. Co. - Greg, 4 men. Set forms to South 1/2 of drive at Coppell Hardware store. Did not install proper width as previously discussed with Scott Balch and the previous foreman for Eagle Const. (Carl) approx. 4 weeks prior. No contact from general contractor to~ sub-contractor prior to starting work. Informed sub-contractor to decrease to proper width and to remove all unstable muck from the grade, caused by the irrigation system. Time: 9:30 AM ±. Contractor departed shortly thereafter to start work on the concrete median nose cap on the north project. Thursday, 9/21/89 North project - Clear, Sunny 62°-85°. SYB Construction Co. - Rick Bennett, 3 men, i backhoe (case' 580-D). Cut and removed bottom section of 14" O.D. steel casing pipe around 8" PVC water main for Skaggs. Cut 8" PVC main and lowered with 8"x24" offset bends and 2 retainer glands. Concrete encased lower section at approx. 4:00 PM. city of Dallas Water Dept. personnel observed and verified the lowering for clearance on the Dallas 30" PCCP. Brooks Maltby - 2 laborers - Installing 3" sweeps and pull boxes at Parkway Blvd. - Denton Tap Road intersection. Proceeding south of Parkway Blvd. with 3" PVC sweeps and pull boxes within the median islands. L.H. Lacy Co. - No work on project Misc. item - city of Coppell street dept. personnel (Per Birdsall) informed me at 3:00 PM that an accident had occurred at the median nose - north side cross over sta. 203+25 ~ last night (approx. 11:30 PM). Woman driving car in northbound lane, drove into and over top of median nose concrete. City crew installed 2 barricades last night at approx. 12:15 AM. Informed Scott Balch at 3:30 PM to remedy any trouble spots within this area through the use of barricades, signs, etc. Note: If left turn lane buttons (Park entrance) had been installed prior to traffic switch, delineation of this area would have been defined, thus guiding traffic away from the opposing median nose. South project - Jim Bowman Const. Co. - Did not show - no work on bridge rails.today. Eagle Const. Co. - Greg and 4 men. Removed unstable soils from area and placed suitable materials back in drive approach right centerline sta. 122+20 (Coppell Hardware Store). Compacted and set forms and tied steel. Inspected at 11:00 AM. Poured concrete at approx. 12:00 noon. Departed project at 2:00 PM ±. L.H. Lacy Co. - Steve, 5 men, 1 laborer, 2 blades. General clean-up of east side of roadway from Braewood to Bethel Road. Median topsoil being placed. Weedeating left side and sta. 135+00 to 140+00 ± (west Parkway along Thweatt property). #2 blade not operating 75% of the time. 2nd clean-up crew appears to be off the project. Barricades and warning signs out of position from Kentucky Fried Chicken sta. 152+50 ± to Dairy Queen sta. 89+00 left of centerline. Friday, 9/22/89 Clear, Sunny 64°-80° winds gusting - North project - Brooks Maltby - 2 laborers - Installing pull. boxes in median islands (2) south of Parkway Blvd. SYB Construction - Randy Bennett, 2 men - Placed additional concrete for thrustblocking on top bends of the offsets, immediately following the installation of City of Dallas 30" waterline. Time placed: 2:00 PM ~. Dallas Water Utilities - Installing shoring system and laid 2-38' joints of 30" PCCP. Misc. Items - Drove project with John Karlsruher and project manager for L.H. Lacy Co. ~(Scott Balch) from 2:45 PM to 3:15 PM. Discussed the need for sandbags to be installed on barricades and signs. Winds have blown down the advance warning signs in transition 192+00 to 206+00 Rt. and divided median nose signs at Parkway Blvd. - Denton Tap Road. Met with Stan Reed, P.E. and John Karlsruher for observation of the Cottonwood Creek Bridge. Time: 4:30 PM to 5:05 PM. South project - Jim Bowman Const. Co. - Did not show. No work on bridge rails today. Eagle Const. Co. - Did not show. No work on 3-30' radii at Braewood (1) and Bethel School Road (2). GTE - Installed waterproof underground splice for Braewood radius (southeast) splice. Compacting ditchline with backhoe. L.H. Lacy Co. - #1 clean-up crew; Steve and 6 men, 1 loader, 1 blade (92 blade down), 1 box blade. Working Steve Parson's frontage and Coppell Medical clinic at Meadowcreek on clean-up. Power broom sweeping east 34' width roadway from Coppell Hardware Store to Grapevine Creek Bridge. Placing topsoil in median islands. Requested Scott Balch to notify barricade company for restoration and re-alignment of 2-way traffic signs, barricades, do not pass signs and speed limit signs. Time: 3:00 PM. Called Scott Balch at 5:10 PM and informed him the barricades and signs have not been taken care of. Saturday, 9/23/89 Clear, Mild, Sunny - Contractor did not work. Sunday, 9/24/89 Clear, Mild, Sunny - Contractor did not work. Barricades have not been serviced as discussed with Scott Balch (L.H. Lacy Co.) on 9/22/89, Friday. -- Monday, 9/25/89 Cool, Mild, 45°-75°. North project - Brooks Maltby, Inc. - 3 men, 1 backhoe. Worked on backhoe from 10:00 AM to 12:00 noon - Flat tire. Installing pull boxes and PVC sweeps on north side of Parkway Blvd.-Denton Tap Road. Lectric Safety Lites - Marvin, 2 men - Installed 8" single reflective yellow buttons in asphalt transition sta. 203+29 to 206+00 at 11:00 'AM to 12:30 PM. Mobilized to south project after lunch. Dallas Water Utilities - Installing 30" PCCP. City of Coppell - Loaded relocated 8" watermain at approx. 11:00 AM for Skaggs/Alpha Beta. Abandoned temporary water bypass at same time. Misc. Items - Drove through project with Ross from L.H. Lacy at 4:00 PM to 4:30 PM i.e.; barricades and transitions. South project - L.H. Lacy Co. - Steve and 7 men, 1 sweeper, 1 blade, 2 small backhoes, 1 laborer, 2'dump trucks. #1 clean-up crew only, 2nd crew departed. Two laborers patching curbs, etc. Two laborers on small backhoes with front end loaders dressing up behind east outside back of curbs. Loader cleaning and loading construction debris from Steve Parson's property Lt. centerline sta. 111+80 to 114+00. Dump trucks hauling to private landowner off Sandy Lake Road behind the Bait Shop. Sweeping roadway surface with power broom. Jim Bowman Const. Co. - James Long, 2 laborers - Mobilized on project in AM and tieing steel for east outside bridge rail to Grapevine Creek Bridge. Marking layout for the forms. Additional anchor belts for bridge lighting arrived at approx. 1:00 PM today. Lectric Safety Lites - 1:00 PM to 4:30 PM - Marvin, 2 men. Installed skip lane buttons from Braewood to Minyards on the inside northbound lane. Suma-Stone - 6 men, 3:00 PM to 5:30 PM. Installed brick pavers to south drive at Coppell Hardware Store. Misc. Items - Drove project with L.H. Lacy representative (Ross) from 4:30 PM to 5:00 PM i.e.; barricades, signs and transitions. Scott Balch informed me at 11:00 AM that they are scheduled for a traffic switch onto the east 34' width roadway (northbound lanes) on Friday, 9/29/89. Concrete barriers are to be installed across the bridge during rail construction. Barricades sandbagged today from 3:00 PM to 4:00 PM. Tuesday, 9/26/89 Clear, Mild, 48°-80° North projeot - Brooks Maltby, Inc. - 3 men, 1 backhoe - In~talling PVC sweeps and pull boxes south of Parkway Blvd. Lectric Safety Lites - Marvin, 2 men and 2 men with sub- contractors. Sandblasting and painting 4"-'white stripe from Parkway Blvd. towards the north to sta. 206+25 ± Rt. centerline (northbound lane) and the southbound lane from 206+25 to 203+29 Lt. centerline. Dallas Water Utilities - Installing 30" PCCP waterline Rt. centerline sta. 168+00. South project - Lectric Safety Lites - Marvin,_.2 men - Installing buttons Rt. centerline sta. 152+17.40 ± to 132+26.90 (Braewood to Bethel School). Layout on buttons Rt. centerline sra. 132+26.90 to 89+00. Jim Bowman Const. Co. - James Long, 5 men - Erecting inside rail wall forms to east outside bridge rail. L.H. Lacy Co. - Steve, 6 men - General clean-up from Joe Bob Fikes business to Dairy Queen Rt. centerline (east 1/2 roadway).